Let X be a random variable having an unknown distribution with mean m=8 and variance s2=9 (standard deviation s =3). Find the following probability: (a) P(-4 <X< 20) (b) P(|X-8| ³ 6)



1
Expert's answer
2021-08-31T17:25:15-0400

By Chebyshev’s Theorem for any numerical data set,

at least "3\/4"  of the data lie within two standard deviations of the mean, that is, in the interval with endpoints "\bar{x}\pm2s" for samples,

at least "8\/9"  of the data lie within three standard deviations of the mean, that is, in the interval with endpoints "\bar{x}\pm3s" for samplesm

at least "1-1\/k^2" of the data lie within "k" standard deviations of the mean, that is, in the interval with endpoints "x\pm ks" for samples, where "k" is any positive whole number that is greater than "1."

Given "\bar{x}=m=8, s=\sqrt{s^2}=3."

(a)


"P(-4<X<20)=P(8-4(3)<X<8+4(3))"

"=1-\dfrac{1}{4^2}=\dfrac{15}{16}"

(b)


"P(|X-8|<6)=P(8-2(3)<X<8+2(3))"

"=\dfrac{8}{9}"